# The Radius Distribution of M dwarf-hosted Planets and its Evolution

### v1

@@ 2.3 Validation
To validate our methodology for reconstructing the planet radius distribution we constructed a sample of synthetic KOIs the same size as the actual one (115 planets, see Sec. 3), drawing radii from the sum of two Gaussian distributions representing super-Earths and sub-Neptunes. The fractional populations in the two distributions were set to 0.4 and 0.6, the centroids were set to 1.3 and 2.1 , and the standard deviations were set to 0.15 and 0.25, respectively. Orbital periods were selected from a power-law distribution between 1 and 40 days that reproduces the observed distribution (Gaidos et al., 2016). The synthetic planets were randomly assigned to the Kepler late K/early M population with measured rotation periods as defined above. To calculate the planet-star radius ratio, the stellar radii were determined using the metallicity-dependent relation of Mann et al. (2015), with metallicities drawn from the Gaussian distribution found among Kepler M dwarfs Gaidos et al. (2016). However for added realism these artificial metallicities were not propagated to the analysis of synthetic KOIs since the metallicities of many KOIs are not well determined; instead, the metallicity-independent relation of Mann et al. (2015) was used. The total probability of detection (including the geometric factor) was calculated for each star-planet pair as described above and the pair accepted conditional to a random uniform draw below this probability. In lieu of actual posterior distributions for the observed radius ratio we adopted Gaussians with fractional standard deviation that are 1/SNR, where SNR is the predicted transit signal-to-noise (Howard et al., 2012):
where is the transit depth (approximately the radius ratio squared), is the measured combined differential photometric precision of the star over 2 hours (Christiansen et al., 2012, , 2 hr is a typical transit duration), the number of transits (equal to the dataspan times the duty cycle divided by the orbital period), and the transit duration (after averaging over a uniformly distributed impact parameter). Host star-planet pairs were generated until the accepted number equaled the size of the actual sample
@@
We ran the synthetic KOIs through the same analysis pipeline as the actual KOIs to produced a simulated observed radius distribution. This is plotted along with the input distribution in Fig. 1. The bimodal distribution is largely reproduced, but with a noticeable broadening of the super-Earth distribution, probably because of the absence of measured [Fe/H] "measurements" and the effect of noise on the radii of the smallest detectable planets. The input ratio of super-Earths to Neptunes (divided by the vertical grey line) is 0.74 while the retrieved ratio is 0.80, an increase of 8%. The increase is due to the greater scatter of the more poorly determined super-Earths into the sub-Neptune range compared to vice-versa. The location of the apparent valley is shifted to larger radii (from 1.64 to 1.74 ) for the same reason.
@@ 3 Results
We refined the radii of 179 KOIs around 119 single late K and early M dwarfs with matches to Gaia DR3 sources and rotation periods in Santos et al. (2019) (Table 1), of which 115 (72 systems) were assigned ages from Gaidos et al. (2023a). The two largest objects in the sample (excluded from further analysis) are the hot Jupiter KOI-254.01 (Kepler-45b, Johnson et al., 2012), and the KOI-2715.01 (Kepler-1321b, Morton et al., 2016) orbiting a late K dwarf (neither shown in Fig. 2). The next largest planet is the sub-Neptune KOI-253.01 (Kepler-2000b, Valizadegan et al., 2023), highlighting how Neptune-size and larger planets are very rare around M dwarfs.
The detection efficiency-corrected radius distribution constructed from the other 113 KOIs (excluding the two largest objects discussed below) is shown by the black line in Fig. 2, with the grey region spanning the 68% confidence interval, and the radii and ages of the KOIs used in this reconstruction are listed in Table 1. The total occurrence is the integral over the distribution and is 1.87 planets per star. The planet radius distribution contains two prominent peaks at about 1.3 (super-Earths) and 2.2 (sub-Neptunes). There is a possible but not statistically significant third peak at 3 , consisting of four planets with a similar size. The super-Earths and sub-Neptunes peaks are separated by an intervening gap or “valley" around 1.6-1.7, as first described for solar-type stars (Owen & Wu, 2013; Fulton et al., 2017) and later tentatively identified around M dwarfs (Hirano et al., 2018; Van Eylen et al., 2021). The stellar mass range of our sample is 0.35-0.67M⊙, with a median of 0.59M⊙, and at the median orbital period of 7.1 days; with these, the mass- and period-dependent relation (Eqn. 11) of Ho & Van Eylen (2023) predicts a radius valley centered at . This is in agreement with the distribution found in this sample (vertical red lines in Fig. 2).
The smallest planets in the sample are KOIs 5692.01 ( ) and 7793.01 ( ), around early M and late-K type dwarfs, respectively. The low detection probabilities of these small objects potentially inflate their occurrence and produce an apparent peak in the radius distribution, but this is not very statistically significant (grey zone in Fig. 2). KOI 314.03 (Kepler-138b), with a radius of , makes a prominent contribution to the distribution at larger radii . The probability of detecting it in the overall sample is very low due to its respectable orbital period (10.3 days), thus it could represent a significant sub-population. However, this object is part of a system of at least three transiting planets, and it was only discovered after the system was analyzed with optimized, automated routines following the discovery of its larger cousins (Batalha et al., 2013).
We subdivided the planets with assigned ages into two nearly equal halves based on the median gyrochronologic age (3.86 Gyr). The sub-Neptune peak is much less prominent in the older sample while the super-Earth is larger, with the ratio between the two (integrated over 1.68–4 and 0.7–1.68, respectively) declining from 2.67 to 0.59 from the younger to older moieties. This is consistent with the evolution of planets by the loss and/or contraction of low molecular-weight envelopes over time. Indeed, the integrated probabilities over mini-Neptunes (1.7–4) fall from 0.65 to 0.31 and is roughly equivalent to the increase in the Earth- and super-Earth (0.7-1.7) from 0.24 to 0.53. There is no evidence for a decrease in the mean radius of the sub-Neptunes with age nor a deepening of the radius valley as has been seen by other analyses of solar-type stars (Petigura et al., 2022; Chen et al., 2022) but the small sample size limits the robustness of this conclusion. Two-sided Kolmogorov-Smirnov and Anderson-Darling tests provide weak statistical support for the young and old samples to be drawn from different populations ( and 0.16, respectively). A plot of radius vs. age (top panel of Fig. 3) shows that while sub-Neptunes decline relative to super-Earths with age, they are still numerous.
We investigated the sensitivity of this result to the choice of divisor age, since the choice of the median is a statistical convenience rather than being physically motivated. Moreover, the gyrochronologic ages have significant uncertainties which can “blur" age dependence. We thus investigated two other divisors; an age of 1.73 Gyr at which the difference between young and old radius distributions is maximized (minimum -value from a two-sided Kolmogorov-Smirnov test) and an age of 2.8 Gyr which falls in a low density region of the age distribution (horizontal dashed and dotted lines in the top panel of Fig. 3). The two sets of alternative distributions are compared in the bottom panel of Fig. 3 and can be compared to that in Fig. 2. All three distributions exhibit an elevated ratio of sub-Neptunes to super-Earths in the younger population: in the case of the median divisor the population evolves from one with greater Neptunes to one elevated in super-Earths and depleted in sub-Neptunes, while in the case of the maximum significance, the population is initially enriched in sub-Neptunes and evolves into one with more equal numbers of sub-Neptunes and super-Earths. In the minimum ambiguity case, the older population becomes enriched in super-Earths.
We also investigated whether the observed radius evolution could be an artefact of variation in stellar properties with age that affect detection555The main sequence lifetime of stars increases with decreasing stellar mass but this is a factor only for host stars more massive than the Sun.. In cool stars, magnetic activity (spots and flaring) is driven by stellar rotation, and hence as stars age and spin down, the noise introduced into lightcurves by such activity decreases (e.g., Davenport, 2016), This means that, all else being equal, smaller and/or more distant planets can be detected around older stars. To evaluate the magnitude of this systematic, we divided our Kepler late K/M dwarf sample with rotation periods in Santos et al. (2019) into 773 (older) stars more slowly rotating than the 4 Gyr M67 gyrochrone of Dungee et al. (2022), and 1936 (younger) stars more rapidly rotating than that gyrochrone; this age division is within error of our median age of 3.86 Gyr. We calculated the ratio of the detection efficiency between the moieties for each planet and plot these vs. planet radius in Fig. 4. All planets in the radius range of interest are slightly less detectable (by 10%) among the more rapidly rotating (“younger") stars than the more slowly rotating (“older") stars and the difference is largest among the smallest planets. However, the differential in detection efficiency between super-Earths and sub-Neptunes is % and cannot explain the factor of 2 difference shown in Fig. 2. (Because we are unable to adequately filter the overall sample for binary stars, which tend to be more rapidly rotating than single stars and also dilute any transit signals, we do not explicitly correct for this difference).
This observed difference in detection efficiency could possibly arise from a statistical difference in host star properties brought about selecting for stars with detectable rotation period. For example, cooler dwarf stars tend to rotate more slowly at a given age (e.g., Curtis et al., 2019) and their rotation is more difficult to detect, due both to the longer time baseline needed to capture variability and the decreasing activity leading to lower spottedness and amplitude of photometric variability. Thus, cooler, smaller stars are more likely to be included in the younger subset and this could bias transit detection statistics, although perhaps in the opposite way as observed. This effect is manifest in the difference in the distribution of the two subsets (Fig. 5a). However, we also find that the distribution of predicted transit signal-to-noise ratio (SNR) of the two subsets is nearly indistinguishable (Fig. 5b); the median values are 0.35 and 0.36. Here, we calculated the relative SNR for a given and as the inverse of the stellar radius times the square root of the brightness times the transit duration, which scales as , where the Gaia magnitude was used for . Cooler dwarf stars are also intrinsically fainter, which apparently cancels out the effect of smaller radii, at least in this field.
Likewise, variation in the planet population with stellar mass combined with a difference in (Fig. 5a), could mimic evolution. An increasing ratio of super-Earths to sub-Neptunes with stellar mass between 0.5 and 1.4M⊙ (Fig. 12 in Petigura et al., 2022), if extending to the M dwarf range, could result in relatively fewer sub-Neptunes around stars assigned older ages. However, the variation is modest (0.04 per 0.1 dex in mass) and in our M dwarf KOI sample there is no trend between host star luminosity and age (Fig. 6) nor planet radius and host-star luminosity (Fig. 7) that would suggest such an effect. Moreover, the general trend is for less massive stars (which would preferentially populate the young subset) to have fewer super
@@
Distributions of planet radius with orbital period and bolometric stellar irradiance or “instellation" are shown in Figs. 8 and 9, respectively. Bolometric luminosities were calculated using interpolations of the corrections to -band luminosities tabulated in Pecaut & Mamajek (2013), and assuming near-circular orbits. We calculated the detection efficiency (i.e. fraction of planets with a given radius and period if distributed uniformly among the late K/early M type stars in our selected target sample using the procedures described in Sec. 2; this is represented as the grey contours in Fig. 8. As expected detection efficiency decreases with smaller planets on wider orbits, which explains the paucity of detections in that domain (lower right corner of Fig. 8).
4 Discussion and Summary
4.1 Features of the M dwarf planet radius distribution
As has been established for planets around solar-type stars (Fulton et al., 2017) and tentatively around cooler stars (Hirano et al., 2018; Van Eylen et al., 2021), the radius distribution of non-Jovian Kepler planets orbiting late K and early M-type stars contains two peaks (super-Earths and sub-Neptunes) with an intervening gap or “valley" at 1.6-1.7. For this sample’s median stellar mass of 0.59M⊙ and median orbital period of 7.1 days, Eqn. 11 in Ho & Van Eylen (2023) predicts a radius valley centered at (vertical red lines in Fig. 2), in agreement with our derived distribution. On the other hand, the radius-period distribution (Fig. 8) suggests that the M dwarf planet radius valley is not markedly dependent on period, contrary to what has been found for solar-type stars (e.g., dashed red line in Fig. 8; Ho & Van Eylen, 2023).
To quantify the slope of a linear boundary (in log-log space) for our K/M dwarf planet sample, we performed a regression of the full 180-KOI sample (including those lacking age assignments) with support vector machines (SVM) and implemented in the Python sci-kit learn package, with choice of regularization parameter , following Van Eylen et al. (2019). Classes were initially assigned simply based on radii smaller (’sub-Neptunes’) or larger (‘super-Earths’) than 1.68, and then iteratively re-classified (10 times) using the SVM. Confidence intervals were calculated with Monte Carlo simulations where planet radii were randomized according to their uncertainties. We find a slope of -0.026 with a 68% confidence range of -0.043 to 0.0, significantly flatter than the -0.09 found by Ho & Van Eylen (2023) for solar-type stars (magenta dashed line in Fig. 10). (Bonfanti et al. (2023) also report a flatter slope of -0.065 for M dwarfs.). Photo-evaporation and core-powered mass loss models predict negative slopes (see Table 6 in Ho & Van Eylen, 2023), while a scenario involving planet formation under gas-depleted conditions predicts a positive slope (Lopez & Rice, 2018).
Cloutier & Menou (2020) found that the planet radius distribution hosted by stars with masses 0.08–0.42M⊙ (corresponding to spectral types M3-M8) differs from that of more massive stars in lacking a significant population of (sub)Neptunes. Ment & Charbonneau (2023) also found a very pronounced dearth of sub-Neptunes around mid-to-late type M dwarfs. Our sample does not substantially overlap in stellar mass with these analyses, but there are no sub-Neptunes detected around the least luminous stars in the sample () and the distribution with luminosity (Fig. 7) is at least consistent with the disappearance of sub-Neptunes at the faint end.
There is a striking paucity of planets significantly smaller than the super-Earth peak at (Fig. 2). Not all of this can be explained by the reduction in detection efficiency (contours in Fig. 8), suggesting that the super-Earth peak is physical. The efficient formation of relatively massive rocky planets around M dwarfs might point to a high density of solids (as “pebbles", planetesimals, or protoplanets) brought about by drift/migration in the inner gas disks around M dwarfs before they are cleared (e.g., Chiang & Laughlin, 2013; Chatterjee & Tan, 2014; Gaidos, 2017).

### v2

@@ 2.3 Validation
To validate our methodology for reconstructing the planet radius distribution we constructed a sample of synthetic KOIs the same size as the actual one (115 planets, see Sec. 3), drawing radii from the sum of two Gaussian distributions representing super-Earths and sub-Neptunes. The fractional populations in the two distributions were set to 0.4 and 0.6, the centroids were set to 1.3 and 2.1 , and the standard deviations were set to 0.15 and 0.25, respectively. These values produces a reasonable radius distribution that roughly approximates the observed distribution but because this a test of the fidelity of recovery, the precise values do not matter. Orbital periods were selected from a power-law distribution between 1 and 40 days that reproduces the observed distribution (Gaidos et al., 2016). Synthetic planets were randomly assigned to the Kepler late K/early M population with measured rotation periods as defined above. To calculate the planet-star radius ratio, the stellar radii were determined using the metallicity-dependent relation of Mann et al. (2015), with metallicities drawn from the Gaussian distribution found among Kepler M dwarfs Gaidos et al. (2016). However, for added realism these artificial metallicities were not propagated to the analysis of synthetic KOIs since the metallicities of many KOIs are not well determined; instead, the metallicity-independent relation of Mann et al. (2015) was used. The total probability of detection (including the geometric factor) was calculated for each star-planet pair as described above and the pair accepted conditional to a random uniform draw below this probability. In lieu of actual posterior distributions for the observed radius ratio we adopted Gaussians with fractional standard deviation that are 1/SNR, where SNR is the predicted transit signal-to-noise (Howard et al., 2012):
@@
where is the transit depth (approximately the radius ratio squared), is the measured combined differential photometric precision of the star over 2 hours (Christiansen et al., 2012, , 2 hr is a typical transit duration), is the number of transits (equal to the dataspan times the duty cycle divided by the orbital period), and the transit duration (after averaging over a uniformly distributed impact parameter). Host star-planet pairs were generated until the accepted number equaled the size of the actual sample We ran the synthetic KOIs through the same analysis pipeline as the actual KOIs to produced a simulated observed radius distribution. This is plotted along with the input distribution in Fig. 1. The bimodal distribution is largely reproduced, but with a noticeable broadening of the super-Earth distribution, probably because of the absence of measured [Fe/H] "measurements" and the effect of noise on the radii of the smallest detectable planets. The input ratio of super-Earths to Neptunes (divided by the vertical grey line) is 0.74 while the retrieved ratio is 0.78, an increase of 5%. The increase is due to the greater scatter of the more poorly determined super-Earths into the sub-Neptune range compared to vice-versa. The location of the apparent valley is shifted to a slightly larger radius (from 1.64 to 1.75 ) for the same reason.
@@ 3 Results
3.1 Planet Radius-Period Distribution
We identified 117 confirmed/validated or candidate planets around 74 single late K and early M dwarfs with matches to Gaia DR3 sources and rotation periods in Santos et al. (2019). Thirteen have NASA Exoplanet Archive dispositions of "candidate" while the rest are "confirmed" (i.e., validated). Orbital periods, radii and ages of KOIs are provided in Table 1. Of these, 114 (in 73 systems) have rotation-based ages: nearly all are from Gaidos et al. (2023) but two missing systems (KOIs 2542 and 2705) were assigned ages using the methods of Gaidos et al. (2023). Four planets in three other systems (KOIs 430, 4957, and 6863) have rotation periods in Santos et al. (2019) that are shorter than the calibrated range in Gaidos et al. (2023) and are not age-dated.
@@ 3.1 Planet Radius-Period Distribution
The two largest objects in the sample, excluded from further analysis, are the hot Jupiter KOI-254.01 (Kepler-45b, Johnson et al., 2012), and the KOI-2715.01 (Kepler-1321b, Morton et al., 2016) orbiting a late K dwarf (neither shown in Fig. 5). The next largest planet is the sub-Neptune KOI-253.01 (Kepler-2000b, Valizadegan et al., 2023), highlighting how Neptune-size and larger planets are very rare around M dwarfs. Figure 2 plots the radius vs. orbital period of the 115 small planets, along with a Gaussian kernel density estimator (KDE) to highlight structure. Planets with assigned ages are plotted as black points; while those lacking ages are plotted in grey.
@@
We calculated the detection efficiency (i.e. fraction of detected planets with a given radius and period if distributed uniformly among the late K/early M type stars), in among 3200-4200K stars in Santos et al. (2019) using the procedures described in Sec. 2. This is represented as the grey contours in Fig. 2. As expected, detection efficiency is lowest for smaller planets on wider orbits, which explains the absence of detections in the lower right corner of the plot. There is a single object, KOI-571.05, which is in a region of particularly low detection efficiency (0.002) and would normally warrant scrutiny as a potential false alarm, but this was validated as a member of the multi-planet Kepler-186 system by Quintana et al. (2014).